# On the Ability of Deep Networks to Learn Symmetries from Data: A Neural Kernel Theory

### v1

@@ 3.3 Multi-layer perceptrons (MLP) on rotated MNIST
Building on the insights from the previous section, we now turn to a more realistic setting. We study the predictions of a MLP in the Neural Tangent Kernel (NTK) limit on partial views of rotated-MNIST, a high-dimensional dataset containing a rotational symmetry.
Datasets and architectures.
@@
We consider the MNIST dataset, and we augment it by means of (discrete) rotations. More precisely, for each MNIST digit, $N$ images are generated by rotation in increments of $2\pi/N$, capturing the full rotational range. This way, we have access to a complete rotational orbit for each sample. This rotation transformation corresponds, up to pixel discretization effects, to the action of a particular representation $g$ of the cyclic group of order $N$ (specifically, the regular representation), acting on the vector space $\mathbb{R}^{28\times 28}$ in which the greyscale MNIST images live. Finally, we normalize the digits to be on the sphere, making their orbits comparable with each other.
@@
We then form pairs of these orbits, each pair consisting of digits from different classes. Given these datasets, we define a regression problem by leaving out one of the points in the two orbits, and asking to find a predictor for this missing sample. Note that unlike in the more realistic scenario presented in Fig. 1, our classes here are constructed from a single seed sample and their orbit, and we only consider two classes. However, we show in Section 3.4 that our conclusions on this simplified setup also apply to the multi-seed-per-class and multi-class problems, with some further approximations.
@@
We now focus on neural networks as our predictors. Specifically, we consider respectively a 1-hidden-layer and a 5-hidden-layer MLPs with ReLU activations (see App. B for more details on the architectures) in the NTK limit. In this limit, the predictions of the networks can be written as the result of Gaussian Process regression, where the kernel function is determined by the architecture of the network. Therefore, we can replicate the analysis performed in the previous section, where we replace the simple Gaussian kernel with a neural kernel. In the following, we refer to our kernel function as $k_{NTK}$, and the respective kernel matrix as $K^{NTK}$. In practice, we use the neural-tangents library (Novak et al.,, 2020) to compute the neural kernels.
@@
Assumption 1: Circularity of the kernel matrix over one orbit.
Firstly, the NTK of a MLP is a dot product kernel. As a consequence, it is not stationary (it is not solely function of the Euclidean distance between pairs of input points); instead, it is a function of the angle between the two input vectors (Neal, 1996b, ). However, in virtue of the simple cyclic nature of our datasets, the kernel matrix $K_{ij}=k(x_{i},x_{j})$ corresponding to the kernel function induced by the MLP on a single class is circulant.
Proof.
Consider class $A$. The cyclic group acts on points of class $A$ linearly via its (matrix) representation $R$; due to the properties of the cyclic group, such representation is orthogonal. The dot product between data points thus only depends on their index difference, as indexed by the group action:
$(x_{i}^{A})^{T}x_{j}^{A}=(R^{i}x_{0}^{A})^{T}(R^{j}x_{0}^{A})=(x_{0}^{A})^{T}R^{-i}R^{j}x_{0}^{A}=(x_{0}^{A})^{T}R^{j-i}x_{0}^{A}$ (8)
As the kernel of a MLP only depends on the scalar product of its inputs, it thus also only depends on the index distance between the data points:
$K_{ij}=k(x_{i}^{A},x_{j}^{A})=k((x_{i}^{A})^{T}x_{j}^{A})=k((x_{0}^{A})^{T}R^{j-i}x_{0}^{A}),$ (9)
and thus the $(i,j)$ entry of the kernel matrix depends only on the difference between the two indices $(i-j)$. The neural kernel computed over a single class is thus circulant. ∎
Assumption 2: Circularity of the kernel matrix over the dataset made of two orbits.
Secondly, MNIST digits from different classes are in general not geometrically interleaved, i.e., a point in one orbit is not equidistant from the “neighboring” points in the other orbit. This can be written as the following dot product condition:
${\color[rgb]{0.23828125,0.39453125,0.6484375}\definecolor[named]{pgfstrokecolor}{rgb}{0.23828125,0.39453125,0.6484375}x_{i}^{A}\cdot x_{i}^{B}}\neq{\color[rgb]{0.4765625,0.57421875,0.7890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.4765625,0.57421875,0.7890625}x_{i}^{A}\cdot x_{i+1}^{B}.}$ (10)
@@
${\color[rgb]{0.94140625,0.3125,0.22265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.3125,0.22265625}x_{i}^{A}\cdot x_{i+1}^{A}}\neq{\color[rgb]{0.9609375,0.59375,0.48828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.59375,0.48828125}x_{i}^{B}\cdot x_{i+1}^{B}.}$ (11)
@@
As a consequence, the following NTK matrix is in general not circulant [^3] We remark, however, that the first diagonal, denoted in black, is constant in virtue of the fact that we normalize each data point to be on the sphere.:
@@
Assumption 3: Symmetrization of the NTK error w.r.t. to class.
Lastly, because of the asymmetry between orbits, removal of a point from one orbit is not in general equivalent to removal of a point from the other orbit. Yet, our formula for the spectral error, operating on the circularized kernel matrix, is intrinsically symmetric. To restore this interchangeability in the standard NTK error, we consider in the following a symmetrized NTK error, which we obtain by averaging the NTK prediction errors that arise by removing a point from either one orbit or the other. It is this symmetrized NTK error that we compare with the spectral error.
@@
Results.
Equipped now with a circulant kernel matrix, we can once again refer to the spectral error formula of Eq. 7 for the estimation of the prediction error associated with performing GP regression under the kernel matrix $\tilde{K}^{NTK}$.
We find that the spectral error tracks the symmetrized NTK error across all sampled pairs (Fig. 4B). This agreement holds across the whole range of explored values for both $\lambda_{N}$ and $\langle\lambda^{-1}\rangle$ (Fig. 4C), and regardless of number of rotation angles (App. D). Moreover, this agreement qualitatively holds for finite-width, trained networks (Fig. 5B and App. E), as well as for deeper MLPs (Fig. 5C and App. E). This suggests that the circularization procedure preserves the overall structure of the problem, and allows us to establish a parallel to the low dimensional case study previously analyzed.
Geometric interpretation.
We can now interpret the quantities involved in the spectral error (Eq. 7) in light of the insights we obtained in the low dimensional case. This affords us a geometric interpretation of the quantities that drive symmetry learning performance. Firstly, we can effectively read the highest frequency eigenvalue $\lambda_{N}$ as a high-dimensional analogue to the orbit separation $\Delta$ in Eq. 5. Indeed, the values of $\lambda_{N}$ and the euclidean distance between orbit averages in input space are, for the explored ranges, in an approximate linear relationship (Fig. 4D). As a consequence (and perhaps tracking intuition), one can see how the separation between the orbits is a key indicator of prediction error. Secondly, we can map the average reciprocal spectrum $\langle\lambda^{-1}\rangle$ to a measure of orbit density, computed as the inverse of the orbit radius [^4] The orbit radius is computed as the distance of any orbit sample to the orbit centroid, in the high-dimensional dataset space. It is a measure of how close the seed image is from the stabilizer of the group action. For example, a seed image that is rotation invariant, such as a perfect ‘o’, is situated on the stabilizer of the rotation group action, leading to a null radius. In our theory, we find that the orbit radii of two orbit-classes to separate is a determinant factor for correct generalization.. We see that these two quantities are in an approximately linear relation (Fig. 4E). These relations enable a geometric interpretation of the quantities that inform class separability, mainly the distance between classes $\Delta$ and the density of orbit-classes. Note, however, that the observed mapping between spectral quantities of the kernel and geometric quantities in input space is empirical and not guaranteed to hold for all datasets or architectures. In particular, we show later that this equivalence breaks for a convolutional neural network with global average pooling at the last layer.
@@ 3.4 Extension to multiple seeds per class and multiple classes
Until now, we have only considered a simplified scenario in which the datasets are composed of a single orbit per class, and where there are only two classes. We now ask whether the theory can also be adapted to a more realistic scenario, where the datasets are composed of multiple seed images per class, and where there are multiple classes, only one of which has a missing angle during training. This setting is the one we presented at the start of this study (Fig 1). Below, we show that our framework can be extended to this scenario.
Extension to multiple seed points per class.
First, we describe how our framework can be extended to multiple seeds per class and two classes. We make a twofold assumption: that (1) the interclass interactions can be effectively summarized by pairwise interactions between orbits of different classes, and (2) orbits belonging to the same class do not interact for the purpose of prediction.
For all pairs of seeds that can be formed across the two classes, we compute a spectral error according to Eq. 7. By doing so, we are considering the case of a missing point in either of these two orbits. We then average the error over all the pairs, obtaining a single, average spectral error.
We compare this average spectral error against the exact NTK error, computed over the entire dataset. We obtain the exact error by running NTK regression on the dataset that contains all orbits of both classes, excluding one angle from all orbits of one class. On said angles, we compute the NTK regression error, averaged over all orbits. We then compute its symmetrized version by swapping the two classes, and averaging the two error values.
We test the agreement between the pair-averaged spectral error and symmetrized NTK error across many trials, where each trial corresponds to a different dataset, obtained by taking a number of seeds ($N_{\mathrm{seed}}=13$) from each of two different, randomly picked MNIST digit classes. We find empirically that the average spectral error correctly predicts exact NTK error across a large number of such trials (Fig. 6A, extended version in App. F).
@@
Figure 6: Application of the spectral theory to a MLP trained on a version of rotated-MNIST comprising multiple seeds per class and multiple classes. A: For datasets comprised of two classes of rotated-MNIST, comprising multiple seeds each ($N_{\mathrm{seed}}=13$), we compare the average spectral error, obtained by averaging over all pairings of orbits in the dataset, with the symmetrized NTK prediction error, computed over all possible missing points. Each dot in the scatter plot represents a different dataset, drawn by randomly selecting two classes from MNIST, and randomly selecting 13 seed images per class. The color coding reflects the percentage of seeds (of both classes) for which the NTK regression gives a correct prediction, understood as agreeing in its sign with the label of the missing points. B: On a multi-class (10 classes of MNIST), multi-seed-per-class ($N_{\mathrm{seed}}=13$) version of rotated-MNIST, we compare the generalization accuracy predicted by our multi-class adapted spectral error, with the one of a normally trained MLP (2 hidden layers, trained with a cross-entropy loss). As the number of points in the orbits increases, both trained and spectral accuracies increase on the classification task, progressively and similarly, suggesting that no mechanism for symmetry learning is present for finite-width trained networks that would be unaccounted for by the spectral theory.
Extension to multiple classes.
@@
We compare the predicted accuracy resulting from this procedure, to the empirical accuracy of a finite-width 2-hidden-layer MLP trained with Adam and a cross-entropy loss (see details of architecture in App. B.3), on a version of rotated-MNIST comprising $N_{\mathrm{seed}}=13$ seeds per class (we limit the number of seeds per class to 13 because of the prohibitive computational cost of computing NTK regression on larger datasets).
The spectral accuracy qualitatively matches the empirical accuracy of the trained network as we vary the number of points composing the orbits (Fig. 6B, extended version in App. Fig. 13). The spectral accuracy curve also mirrors the empirical curves obtained by training various architectures (MLP, CNN, ViT) on a full version of rotated-MNIST at the beginning of this study (Fig. 1).
In conclusion, our spectral theory, which only considers pairwise interactions between seeds across classes, and lends itself to a simple geometric interpretation of the factors leading to correct generalization (namely class distance and orbit density), qualitatively recapitulates the lack of generalization of conventionally trained finite-width networks on a realistic dataset, rotated-MNIST, comprising multiple classes and multiple seeds per class. This result provides strong evidence that the ability of conventional deep networks to generalize on a symmetric dataset is essentially dictated by the local geometric properties of the dataset, and that no specific mechanism allows such networks to “grok” (learn from examples) the non-local, symmetric structure of a prediction problem.

### v2

@@ 4.4 Multi-layer Perceptrons on Rotated-MNIST
Building on the insights from the previous section, we now turn to a more realistic setting. We study the predictions of Multi-layer Perceptrons (MLP) in the Neural Tangent Kernel limit (NTK), on partial views of rotated-MNIST, a high-dimensional dataset containing a rotational symmetry.
Datasets and architectures
@@
We consider the MNIST dataset, and we augment it by means of (discrete) rotations. More precisely, for each MNIST digit, $N$ images are generated by rotation in increments of $2\pi/N$, capturing the full rotational range. This way, we have access to a complete rotational orbit for each sample. This rotation transformation corresponds, up to pixel discretization effects, to the action of a particular representation $g$ of the cyclic group of order $N$ (specifically, the regular representation), acting on the vector space $\mathbb{R}^{28\times 28}$ in which the greyscale MNIST images live. Additionally, we normalize the digits to be on the sphere, making their orbits comparable with each other.
@@
We then create datasets by forming pairs of these orbits, each pair consisting of digits from different classes. Given these datasets, we define a regression problem by leaving out one of the points in the two orbits, and asking to find a predictor for this missing sample. Note that unlike in the more realistic scenario presented in Fig. 1, our classes here are constructed from a single seed sample and their orbit, and we only consider two classes. However, we will show in Section 4.5 that our conclusions on this simplified setup also apply to the multi-seed-per-class and multi-class problems, under further approximations.
@@
We now focus on neural networks as our predictors. Specifically, we consider respectively a 1-hidden-layer and a 5-hidden-layer MLPs with ReLU activations (see App. D for more details on the architectures) in the NTK limit. In this limit, the predictions of the networks can be written as the result of kernel regression, where the kernel function is determined by the architecture of the network. Therefore, we can replicate the analysis performed in the previous section, where we replace the simple Gaussian kernel with a neural kernel. In the following, we refer to our kernel function as $k_{NTK}$, and the respective kernel matrix as $K^{NTK}$. In practice, we use the neural-tangents library (Novak et al., 2020) to compute the neural kernels.
@@
Firstly, unlike the RBF kernel, the MLP kernel is not stationary, so Lemma 5 does not apply to the MLP kernel. However, we can show that the MLP kernel, in virtue of being a dot product kernel (Neal, 1996), also produces a circulant Gram matrix over an orbit of the cyclic group:
Lemma 6
The Gram matrix of a dot product kernel is circulant over a single orbit of the cyclic group.
Proof  The cyclic group acts on a seed point linearly via its (matrix) representation $R$; due to the properties of the cyclic group, such representation is orthogonal. The dot product between data points thus only depends on their index difference, as indexed by the group action:
$(x_{i}^{A})^{T}x_{j}^{A}=(R^{i}x_{0}^{A})^{T}(R^{j}x_{0}^{A})=(x_{0}^{A})^{T}R^{-i}R^{j}x_{0}^{A}=(x_{0}^{A})^{T}R^{j-i}x_{0}^{A}$
As the kernel of a dot product only depends on the scalar product of its inputs, it thus also only depends on the index distance between the data points:
$K_{ij}=k(x_{i}^{A},x_{j}^{A})=k((x_{i}^{A})^{T}x_{j}^{A})=k((x_{0}^{A})^{T}R^{j-i}x_{0}^{A}),$
and thus the $(i,j)$ entry of the Gram matrix depends only on the difference between the two indices $(i-j)$, modulo $N$ because of the properties of the rotation matrix. The Gram computed over a single orbit is thus circulant.
Next, we remind the reader that the datasets we consider here are composed of two orbits, obtained by applying the cyclic group action to two different MNIST digits. Unlike in the synthetic dataset of the previous section, these two orbits are not geometrically interleaved, as we have no control over the position of the seed points given by the MNIST digits. This leads us to make the following approximation and adjustment.
Approximation: Circularity of the kernel matrix over the dataset made of two orbits
MNIST digits from different classes are in general not geometrically interleaved, i.e., a point in one orbit is not equidistant from the “neighboring” points in the other orbit. This can be written as the following dot product condition:
${\color[rgb]{0.23828125,0.39453125,0.6484375}\definecolor[named]{pgfstrokecolor}{rgb}{0.23828125,0.39453125,0.6484375}x_{i}^{A}\cdot x_{i}^{B}}\neq{\color[rgb]{0.4765625,0.57421875,0.7890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.4765625,0.57421875,0.7890625}x_{i}^{A}\cdot x_{i+1}^{B}.}$
@@
${\color[rgb]{0.94140625,0.3125,0.22265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.3125,0.22265625}x_{i}^{A}\cdot x_{i+1}^{A}}\neq{\color[rgb]{0.9609375,0.59375,0.48828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.59375,0.48828125}x_{i}^{B}\cdot x_{i+1}^{B}.}$
@@
As a consequence, the following NTK matrix is in general not circulant: [^6] We can note that the first diagonal (denoted in black) is constant in virtue of the fact that we normalize each data point to be on the sphere.
@@
Figure 4: Analysis of the prediction behavior of a MLP on pairs of orbits from rotated-MNIST. A: We take samples of two different MNIST digits and generate their rotation orbits. The task is to predict the label value of a leave-out point in one of the two orbits. We use angle steps of 45 degrees, amounting to $N_{rot}=8$. We show the points in a reduced-dimensional space obtained by performing PCA on the dataset. B: Scatter plot of the NTK error computed in the standard way (exact) against our spectral error (Eq. 1). Each dot corresponds to a different dataset, obtained by randomly picking pairs of digits of different classes. The color coding corresponds to the number of classification errors incurred by the symmetrized NTK regression, obtained by excluding a point from class A, then from class B, and counting the number of classification errors (0, 1 or 2), understood as a disagreement in sign between NTK prediction and label of the missing point. C: Comparison between the spectral and exact NTK error across different values of $\lambda^{-1}_{N}$ and $\langle\lambda^{-1}\rangle$. D: Comparison between the values of $\lambda_{N}$ and $\Delta^{2}$. The former is the highest frequency component of the neural kernel matrix (inverse of the numerator in Eq. 1), while the latter is the distance in pixel space between the averages (centroids) of the two orbits. E: Comparison between the values of $\langle\lambda^{-1}\rangle$ and $1/\rho$ (see main text).
@@
Adjustment: Symmetrization of the NTK error w.r.t. to class
Because of the asymmetry between orbits, removal of a point from one orbit is not in general equivalent to removal of a point from the other orbit. Yet, our formula for the spectral error, operating on the circularized kernel matrix, is intrinsically symmetric. To restore this interchangeability in the standard NTK error, we consider in the following a symmetrized NTK error, which we obtain by averaging the NTK prediction errors that arise by removing a point from either one orbit or the other. It is this symmetrized NTK error that we compare with our spectral error.
Results
Equipped now with a circulant kernel matrix, we can once again refer to the spectral error formula of Eq. 1 for the estimation of the prediction error associated with performing kernel regression under the kernel matrix $\tilde{K}^{NTK}$.
We find that the spectral error tracks the symmetrized NTK error across all sampled pairs (Fig. 4A-B). This agreement holds across the whole range of explored values for both $\lambda_{N}$ and $\langle\lambda^{-1}\rangle$ (Fig. 4C), and regardless of the number of rotation angles (App. F). Moreover, this agreement qualitatively holds for finite-width, trained networks (Fig. 5B and App. G), as well as for deeper (5-layer) MLPs (Fig. 5C and App. G). This suggests that the circularization procedure preserves the overall structure of the problem, and allows us to study generalization through a geometric lens as in the previous section.
@@ Results
Geometric interpretation
The quantities involved in the spectral error (Eq. 1) are computed in kernel space, not in input (i.e., dataset) space, which makes their interpretation difficult. However, for simple kernels such as the MLP kernel, we will now show empirically that these quantities correlate with geometric quantities in input space, allowing us to understand the geometric factors underlying generalization.
Firstly, the highest frequency eigenvalue $\lambda_{N}$ correlates with the orbit separation in input space $\Delta$ (Eq. 2). Indeed, the values of $\lambda_{N}$ and the euclidean distance between orbit averages in input space are, for the explored ranges, in an approximate linear relationship (Fig. 4D). As a consequence, one can see how the separation between the orbits is a key indicator of prediction error for the MLP kernel.
Secondly, we can map the average reciprocal spectrum $\langle\lambda^{-1}\rangle$ to a measure of orbit density in input space, computed as the inverse of the orbit radius. The orbit radius is computed as the distance of any orbit sample to the orbit centroid, in the high-dimensional dataset space. [^7] The orbit radius is a measure of how close the seed image is from the stabilizer of the group action. For example, a seed image that is rotation invariant, such as a perfect ‘o’, is situated on the stabilizer of the rotation group action, leading to a null radius. We see that these two quantities are in an approximately linear relation (Fig. 4E).
These relations enable a geometric interpretation of the quantities that determine generalization success: generalization success depends on the distance between classes $\Delta$ and the density of orbits. Note, however, that the observed mapping between spectral quantities of the kernel and geometric quantities in input space is empirical and not guaranteed to hold for all architectures. In particular, we show later that this equivalence breaks for a convolutional neural network with global average pooling at the last layer, making the interpretation of the spectral formula terms less direct.
@@ 4.5 Extension to Multiple Seeds per Class and Multiple Classes
Until now, we have only considered a simplified scenario in which the datasets are composed of a single orbit per class, and where there are only two classes. We now ask whether the theory can also be adapted to more a realistic scenario, where datasets are composed of multiple seed images per class, and where there are multiple classes, only one of which has a missing angle during training. This setting is the one we presented at the start of this study (Fig 1). Below, we show that our framework can be extended to this scenario.
Extension to multiple seed points per class
First, we describe how our framework can be extended to multiple seeds per class and two classes. Since our spectral formula only applies to datasets made of two orbits (one for each class), we make the simplifying assumption that pairs of orbits interact linearly to predict the average error on a missing point. In other words, we average the spectral error computed over each possible pairs of orbits taken across classes.
Specifically, for all pairs of seeds that can be formed across the two classes, we compute a spectral error according to Eq. 1. By doing so, we are considering the case of a missing point in either of these two orbits. We then average the error over all the pairs, obtaining a single, average spectral error.
We compare this average spectral error against the exact NTK error, computed over the entire dataset. We obtain the exact error by running NTK regression on the dataset that contains all orbits of both classes, excluding one angle from all orbits of one class. On said angles, we compute the NTK regression error, and average it over missing points from all orbits of that class. We then compute its symmetrized version by swapping the role of the two classes, and averaging the two error values.
We test the agreement between the pair-averaged spectral error and symmetrized NTK error across many trials, where each trial corresponds to a different dataset, obtained by taking a number of seeds ($N_{\mathrm{seed}}=13$) from each of two different, randomly picked MNIST digit classes. We limit the datasets to contain 13 seeds per class for computational reasons. We find empirically that the average spectral error correctly predicts exact NTK error across a large number of such trials (Fig. 6A, extended version in App. H). We note that the spectral error does not capture well the magnitude of the exact error anymore (in reason of the additional assumptions needed in the multi-seed case). However, the spectral error still correlates well with the exact error across datasets.
@@
Figure 6: Application of the spectral theory to a MLP trained on a subsection of rotated-MNIST comprising multiple seeds per class and multiple classes. A: For datasets comprised of two classes of rotated-MNIST, comprising multiple seeds each ($N_{\mathrm{seed}}=13$), we compare the average spectral error, obtained by averaging over all pairings of orbits in the dataset, with the symmetrized NTK prediction error, computed over all possible missing points. Each dot in the scatter plot represents a different dataset, drawn by randomly selecting two classes from MNIST, and randomly selecting 13 seed images per class. The color coding reflects the percentage of seeds (of both classes) for which the NTK regression gives a correct prediction, understood as agreeing in its sign with the label of the missing points. B: On a multi-class (10 classes of MNIST), multi-seed-per-class ($N_{\mathrm{seed}}=13$) version of rotated-MNIST, we compare the generalization accuracy predicted by our multi-class adapted spectral error, with the one of a normally trained MLP (2 hidden layers, trained with a cross-entropy loss). As the number of points in the orbits increases, both trained and spectral accuracies increase on the classification task, progressively and similarly, suggesting that no mechanism for symmetry learning is present for finite-width trained networks that would be unaccounted for by the spectral theory.
Extension to multiple classes
@@
We compare the predicted accuracy resulting from this procedure, to the empirical accuracy of a finite-width 2-hidden-layer MLP trained with Adam and a cross-entropy loss (see details of architecture in App. D.3), on a version of rotated-MNIST comprising $N_{\mathrm{seed}}=13$ seeds per class (we limit the number of seeds per class to 13 because of the prohibitive computational cost of computing NTK regression on larger datasets).
The spectral accuracy qualitatively matches the empirical accuracy of the trained network as we vary the number of points composing the orbits (Fig. 6B, extended version in App. Fig. 14). The spectral accuracy curve also mirrors the empirical curves obtained by training various architectures (MLP, CNN, ViT) on a full version of rotated-MNIST at the beginning of this study (Fig. 1). We note that classification accuracy increases with the number of sampled angles until it saturates to perfect classification accuracy. The saturation effect comes from the fact that we evaluate the network on a classification task. When the elements of the orbits are dense enough in kernel space, the error in prediction becomes small enough that the test point is systematically correctly classified.
In conclusion, our spectral theory, which only considers pairwise interactions between seeds across classes, and lends itself to a simple geometric interpretation of the factors leading to correct generalization (namely class distance and orbit density), qualitatively recapitulates the lack of generalization of conventionally trained finite-width networks on a realistic dataset, rotated-MNIST, comprising multiple classes and multiple seeds per class. Moreover, both theory and experiments show that the error is a progressive function of these two quantities, and that there isn’t a phase transition where the network would suddenly learn a function that captures the symmetry of the problem. Taken together, these results provide strong evidence that the ability of conventional deep networks to generalize on symmetric datasets is essentially dictated by local geometrical properties of the datasets, and that no specific mechanism exists that would allow deep networks to learn from examples the non-local, symmetric structure of a prediction problem.